Opened 10 years ago

Closed 9 years ago

Last modified 3 years ago

#266 closed flight processing (fixed)

GB08/08, flight day 268b/2009, Wytham woods

Reported by: mark1 Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_268b_Wytham_woods

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

Priority: alpha-4-High

PI: Richard Harding

Sensors:

  • Eagle (09/04/2010)
  • Hawk (09/04/2010)
  • Leica LIDAR (25/03/2010)

Change History (25)

comment:1 Changed 10 years ago by mark1

Still awaiting flight logsheet.

Also, errors when generating DEM in Grass but it still output a DEM, check it is good when processing.

When logsheet arrives check eagle / hawk files and rename them - currently all named with 268c - need to check they are for 268b and not 268c.

comment:2 Changed 10 years ago by mark1

Eagle and Hawk files in wrong project.

Eagle and Hawk files have been moved to GB08_10-2009_268c_Denny_wood/. Eagle files from boresight_2009-268a_little_riss have moved into this project. Currently no Hawk files.

comment:3 Changed 10 years ago by mark1

  • Milestone set to 2009 Data processing completion

comment:4 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:5 Changed 10 years ago by mggr

Hawk data recovered from aircraft disks, put in directory and renamed from Hawk268b-09-* to SWIR268b-09-*.

comment:6 Changed 10 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-3.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-5.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-6.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_08-2009_268b_Wytham_woods/hawk/SWIR268b-09-7.hdr: predicted length half recorded length, frame rate doubling problem?

comment:7 Changed 10 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

Starting LIDAR processing

comment:8 Changed 10 years ago by emca

LIDAR delivery directory created at
~airborne/workspace/GB08_08-2009_268b_Wytham_woods/delivery/20100322
Ready for delivery check.

Starting Hyperspectral processing

comment:9 Changed 10 years ago by mark1

LIDAR delivery checked. Data viewed in lag and looks good.
Ready to deliver.

comment:10 Changed 10 years ago by mark1

  • Description modified (diff)

LIDAR dispatched to PI on 25/03/2010

comment:11 Changed 10 years ago by emca

No specim sync for Eagle line 1 or any of the Hawk lines.

Per line SCT's

Line Eagle Hawk
1 -0.01 -0.02
2 0.04 -0.04
3 -0.02 0.08
4 -0.03 -0.02
5 -0.04 -0.03
6 -0.02 -0.01
7 0.02 -0.02

comment:12 Changed 10 years ago by emca

Hyperspectral data ready for delivery check at
~airborne/workspace/GB08_08-2009_268b_Wytham_woods/delivery/20100331

comment:13 Changed 10 years ago by knpa

Delivery Check Complete.

All looks good apart from the following:

  1. One misnamed file: e268061b.bil should be e268b061b.bil
  2. File e268061b.bil will not open - unix says it is a "DOS executable (device driver)"
  3. Need hawk and eagle mosaic's overlayed with vectors screenshots
  4. No listing for bin dir in Readme contents
  5. README_DEM_MAY_NOT_BE_ACCURATE.txt: needs to be unix2dos'd
  6. Example azexhdf command is wrong:

azexhdf -h e268b011b.hdf -G e268b013b.hdf
Should be something like:
azexhdf -h e268b013b.hdf -G e268b013b.tif

comment:14 Changed 10 years ago by emca

File name amended and read_me corrected
Hyperspectral data sent to PI 09/04/2010

comment:15 Changed 10 years ago by emca

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

LIDAR and Hyperspectral data rsync'd to
arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_268b_Wytham_woods

Ready for archiving

comment:16 Changed 9 years ago by knpa

  • Owner changed from emca to knpa

Beginning archiving.

comment:17 Changed 9 years ago by knpa

leica/logs directory missing

.reg file missing

comment:18 Changed 9 years ago by knpa

Vector screenshots not present in lidar or hyperspectral

comment:19 Changed 9 years ago by knpa

Level 3s missing.

comment:20 Changed 9 years ago by knpa

Finished preparing for archiving, ready to be downloaded by NEODC.

comment:21 Changed 9 years ago by knpa

Transferred to, and backed-up by, NEODC.

Moved to archive disk.

comment:22 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace version.

Closing ticket.

comment:23 Changed 7 years ago by besm

Delivery checked reprocessed hyperspectral. Looks fine.

comment:24 Changed 6 years ago by tipo

Reprocessed nav and hyperspectral.

comment:25 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.